# DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures

### v1

@@ A.4.8 Performance of DeepSeek-V3
| Geom Graph* | Construct | 0.04 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
| Network | Bloom Filter* | Compound | 0.00 (0.00) | 0.07 (0.00) | 0.00 (0.00) |
@@ A.5 Accuracy by prompting methods across instruction-tuned models
This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None. The results are shown in Table 20 (GPT-4.1), Table 21 (Gemini-2.0-Flash), Table 22 (Claude-3.5-Sonnet), Table 23 (DeepSeek-V3), and Table 24 (Llama-3.3).
@@
A.6 DSR-Bench-spatial supplementary materials
Figure 3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn. These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section 4.2.
A.7 DSR-Bench-natural supplementary materials
Figure 4 illustrates our generation process for natural language prompts in DSR-Bench-natural. For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in A.7). This narrative is then paraphrased into five variants using GPT-4o. For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in A.7). In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
@@ A.7 DSR-Bench-natural supplementary materials
We construct a scenario in which a clinic uses a BST to store patient appointments, ordered by appointment time and tie-broken by the patient’s name. The insert operation adds a name, time) appointment to the tree, while the delete operation corresponds to a patient canceling their appointment. To retrieve all records, a pre-order traversal of the tree is performed.
@@
A.8 Levenshtein distance: an auxiliary metric for DSR-Bench
@@
In addition to the binary (0/1) accuracy reported in the main text, DSR,-Bench includes an optional evaluation metric based on Levenshtein distance, which measures the minimum number of single-character insertions, deletions, or substitutions needed to transform one string into another. As a continuous metric, it captures degrees of error that binary accuracy flattens. For instance, given the correct output [1,,3,,6], the prediction [3,,1,,6] is clearly closer than [0,,0,,0], and Levenshtein distance reflects that nuance.
@@
However, this granularity can also blur important semantic distinctions. A syntactically well-formed but semantically incorrect output may still receive a high score, especially when the expected answer is long or formatted. When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein. For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table 25).

### v2

@@ A.4.9 Performance of DeepSeek-V3
| Geom Graph* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
| Hybrid | Bloom Filter* | Compound | 0.00 (0.00) | 0.07 (0.00) | 0.00 (0.00) |
@@ A.5 Accuracy by prompting methods across instruction-tuned models
This section presents additional accuracy tables for tasks in DSR-Bench, evaluating each instruction-tuned model across five prompting methods: Stepwise, 0-CoT, CoT, 3-shot, and None. The results are shown in Table˜23 (GPT-4.1), Table˜24 (Gemini-2.0-Flash), Table˜25 (Claude-3.5-Sonnet), Table˜26 (DeepSeek-V3), and Table˜27 (Llama-3.3).
@@
A.6 Additional analysis on CoT prompting
In Section˜4.1.1, we analyzed various prompting strategies with instruction-tuned models. We further examined the CoT method by inspecting its reasoning outputs, which offers insights we hope will benefit practitioners and researchers:
CoT offers limited benefits for well-known tasks. For familiar problems that models are likely to have encountered during pretraining (such as array, queue, and Binary Search Tree), CoT prompting yields only marginal gains. This suggests that models already possess internalized procedures for these tasks and can execute them reliably without additional reasoning steps. Simpler prompts can be more cost efficient and effective.
Without careful design, CoT can hurt performance. In more complex scenarios, such as the hashmap compound task, using CoT with structured JSON-style reasoning actually degraded accuracy. When the reasoning format was changed to natural language, performance recovered to match the baseline (0-CoT). This highlights that the effectiveness of CoT is highly sensitive to prompt design, and poorly chosen reasoning formats may introduce unnecessary complexity rather than aiding problem solving.
A.7 The spatial suite supplementary materials
Figure˜3 presents example illustrations of the non-uniform input distributions: circles, moons, and blobs, which were adopted from scikit-learn. These synthetic patterns are used to evaluate whether models can adapt to irregular and non-uniform spatial distributions, an essential aspect of real-world data, as discussed in Section˜4.2.
We observe a performance drop in KD-Tree tasks when the input data is non-uniform. At the problem level, uniform and non-uniform KD-Tree tasks appear equally difficult for a human reasoner. Across 30 questions per group, duplicated indices, which trigger tie-breaking, occur at similar rates on both axes (x: 32 vs. 33; y: 26 vs. 28), and the total number of median operations is comparable (156 vs. 145). This suggests the surface-level difficulty is well matched across distributions. A closer inspection of errors reveals two main causes:
Non-uniform data often increases ambiguity near the global median.
Non-uniform distributions often produce clusters of similar or identical coordinates near the global median, making it harder for the model to select the correct split. Since the root node is determined by a median over the entire dataset, this ambiguity leads to more errors. Supporting this, we observed 13 out of 30 correct root nodes on uniform data, but only 5 out of 30 on circle (non-uniform) data.
Even simple tasks, such as median calculation, can be challenging for LLMs.
On non-uniform data, models exhibited 28.6% more cases of violating the prompt’s “latter-median” rule by instead choosing the “former-median,” and 50% more instances of selecting a completely incorrect median (neither the former nor latter). Additionally, while no axis confusion errors occurred on uniform data, the model made three such errors on non-uniform data, mistakenly selecting the wrong axis before performing median calculation.
A.8 The natural suite supplementary materials
Figure˜4 illustrates our generation process for natural language prompts in DSR-Bench-natural. For each data structure task, we begin by manually writing an initial scenario narrative (see On a sunny afternoon… arrived in the queue” in Section˜A.8). This narrative is then paraphrased into five variants using GPT-4o. For each supported operation type (e.g., enqueue, dequeue), we generate 30 paraphrased operation templates in a similar manner (e.g., With money in hand, Yuki Lopez stepped into the queue.” corresponds to “enqueue Yuki Lopez” in Section˜A.8). In total, each task is backed by a pool of five scenario descriptions and 30 templates per operation type.
@@ A.8 The natural suite supplementary materials
We construct a scenario in which a clinic uses a BST to store patient appointments, ordered by appointment time and tie-broken by the patient’s name. The insert operation adds a (name, time) appointment to the tree, while the delete operation corresponds to a patient canceling their appointment. To retrieve all records, a pre-order traversal of the tree is performed.
@@
A.9 The code suite supplementary materials
Our benchmark is designed to evaluate general reasoning in LLMs without relying on external tools such as code interpreters or formal solvers. The focus is on the underlying reasoning ability that drives broad problem-solving, beyond coding or domains where code is directly applicable. This follows recent evaluation practices, such as Gemini-Deep-Think and OpenAI’s IMO assessments (Luong & Lockhart, 2025; Wei, 2025), which deliberately disallowed code or tool use (e.g., Lean) to test end-to-end reasoning. By using data structures as controlled, interpretable settings for structural reasoning—a core component of general reasoning—DSR-Bench differs from code-synthesis benchmarks that emphasise syntax or API knowledge and are often vulnerable to data leakage.
To additionally probe how code generation may support structural reasoning, we conducted an ablation across six models, including reasoning models with competitive performance on code-synthesis benchmarks. We cover seven data structures of varying difficulty (Array, Queue, Hashmap, Heap, DAWG, Geom Graph, Graph-Natural) and average results over three runs. Three code-generation modes were tested, with prompts shown below:
CodeOnly: “Your answer should be a Python function ‘def solution()’ that takes no inputs, solves the problem, and returns the final solution in the expected format. Output the code only.”
CodeEnforce: “You should write code to solve the problem, then reason through its execution to explain what the output would be. Your final answer should be the output itself.”
CodeMaybe: “You can write code to help solve the problem, or solve it directly. If you use code, reason through its execution to explain what the output would be. Your final answer should be the output itself.”
| GPT-4.1 | o4-mini | |||||||
| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | None | CodeMaybe | CodeEnforce | CodeOnly |
| Array | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
| Queue | 0.82 | 0.84 | 0.89 | 0.98 | 1.00 | 1.00 | 1.00 | 1.00 |
| Hashmap | 0.19 | 0.11 | 0.09 | 1.00 | 0.89 | 0.89 | 0.86 | 0.94 |
| Heap | 0.58 | 0.53 | 0.53 | 1.00 | 0.44 | 0.73 | 0.70 | 0.53 |
| DAWG | 0.16 | 0.17 | 0.11 | 0.90 | 0.49 | 0.46 | 0.30 | 0.56 |
| Geom Graph | 0.03 | 0.02 | 0.04 | 0.93 | 0.83 | 0.98 | 0.97 | 0.99 |
| Graph-Natural | 0.01 | 0.00 | 0.00 | 0.86 | 0.43 | 0.26 | 0.39 | 0.69 |
| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | Accuracy | Frequency |
| Array | 1.00 | 1.00 | 1.00 | 0.01 | 1.00 | 0.01 |
| Queue | 0.87 | 0.86 | 0.87 | 0.13 | 0.93 | 0.14 |
| Hashmap | 0.28 | 0.49 | 0.49 | 0.62 | 0.93 | 0.67 |
| Heap | 0.23 | 0.28 | 0.34 | 0.83 | 1.00 | 0.83 |
| DAWG | 0.18 | 0.20 | 0.17 | 0.00 | 0.00 | 0.01 |
| Geom Graph | 0.07 | 0.03 | 0.03 | 0.68 | 0.76 | 0.90 |
| Graph-Natural | 0.00 | 0.00 | 0.00 | 0.82 | 0.82 | 1.00 |
| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | Accuracy | Frequency |
| Array | 1.00 | 1.00 | 1.00 | 0.83 | 1.00 | 0.83 |
| Queue | 1.00 | 1.00 | 1.00 | 0.99 | 1.00 | 0.99 |
| Hashmap | 0.58 | 0.17 | 0.13 | 0.57 | 0.95 | 0.60 |
| Heap | 0.36 | 0.37 | 0.50 | 0.37 | 0.42 | 0.87 |
| DAWG | 0.61 | 0.47 | 0.45 | 0.33 | 0.90 | 0.37 |
| Geom Graph | 0.19 | 0.77 | 0.53 | 0.64 | 0.98 | 0.65 |
| Graph-Natural | 0.12 | 0.10 | 0.09 | 0.24 | 0.56 | 0.43 |
| Claude-3.5-Sonnet | Claude-3.7-Sonnet | |||||||
| Structure | None | CodeMaybe | CodeEnforce | CodeOnly | None | CodeMaybe | CodeEnforce | CodeOnly |
| Array | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
| Queue | 0.87 | 0.86 | 0.84 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
| Hashmap | 0.37 | 0.27 | 0.27 | 1.00 | 0.71 | 0.73 | 0.65 | 1.00 |
| Heap | 0.53 | 0.63 | 0.63 | 0.93 | 0.89 | 0.85 | 0.82 | 0.98 |
| DAWG | 0.20 | 0.17 | 0.13 | 0.20 | 0.17 | 0.18 | 0.15 | 0.89 |
| Geom Graph | 0.10 | 0.09 | 0.10 | 0.98 | 0.04 | 0.23 | 0.20 | 0.96 |
| Graph-Natural | 0.00 | 0.00 | 0.00 | 0.10 | 0.01 | 0.02 | 0.01 | 0.23 |
While GPT-4.1 and o4-mini generally perform well, we observe that Claude models also perform competitively, particularly in spatial data structure tasks such as Geom Graph. The only exception is Graph-Natural, suggesting that Claude models struggle with handling natural language ambiguity. On the other hand, Gemini models tend to return final answers directly rather than code in CodeOnly mode, despite explicit instructions (“Your answer should be a Python function… Output the code only.”). This indicates weaker instruction-following in code generation compared to other models and contributes to their low performance. We investigated this further and presented a new table showing (i) task accuracy when code is written (Code Accuracy) and (ii) the code-writing frequency of Gemini models (Code Frequency). We observe that Gemini models also perform on par when code is written.
Models cannot reason over the code they write.
Performance in CodeMaybe and CodeEnforce is on par with our original setup, indicating that writing code does not improve reasoning when models must internally simulate it. This reinforces our central claim: LLMs still struggle with structural reasoning, even when guided by their own code.
Code helps when tasks align with memorized patterns.
In CodeOnly, models perform well on Geom Graph (k-dimensional graphs embedded in geometric space, which is a standard data structure widely used in computer graphics), whose code implementation is more available online. However, they struggle with the less familiar DAWG” (directed acyclic word graph), where we define it with customized constraints to enforce a unique output. This suggests that performance may reflect memorization rather than true reasoning.
CodeOnly falters on natural language task variants.
In Graph-Natural, models rely on brittle pattern matching (e.g., mapping “A space tunnel links planet1 and planet2” to “G.add_edge(planet1, planet2)”), but often fail to cover all phrasing variations or misinterpret descriptions (e.g., missing “Couriers frequently travel the tunnel that connects planet1 to planet2”). This highlights a key limitation that models still struggle to apply structural reasoning to ambiguous natural language scenarios, even with external tools.
Ablation with ChatGPT web UI.
Our benchmark was evaluated via APIs. In contrast, the ChatGPT web UI has a built-in code interpreter that executes code automatically when needed. We tested o4-mini on ChatGPT with DAWG+CodeMaybe. The model scored 0.70, generating and executing code in 20 out of 30 cases. We observed that the model typically uses code for complex algorithmic components, while relying on natural language reasoning for the rest.
A.10 Levenshtein distance: an auxiliary metric for DSR-Bench
@@
In addition to the binary (0/1) accuracy reported in the main text, DSR-Bench includes an optional evaluation metric based on Levenshtein distance, which measures the minimum number of single-character insertions, deletions, or substitutions needed to transform one string into another. As a continuous metric, it captures degrees of error that binary accuracy flattens. For instance, given the correct output [1,3,6], the prediction [3,1,6] is clearly closer than [0,0,0], and Levenshtein distance reflects that nuance.
@@
However, this granularity can also blur important semantic distinctions. A syntactically well-formed but semantically incorrect output may still receive a high score, especially when the expected answer is long or formatted. When averaged over 30 test cases, models with large gaps in binary accuracy can appear deceptively similar under Levenshtein. For example, the Skip List and DSU compound tasks yield the same Levenshtein score (0.75), despite the former achieving more than triple the binary accuracy (Table˜32).
@@
A.11 Failure rates of JSON parsing via Structured Output
| Model | Array | Priority Queue | Hashmap | RB Tree | Geom Graph | Bloom Filter |
| Llama3.3 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| GPT-4.1 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| DeepSeek-Chat | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| DeepSeek-R1 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| o4-mini | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| Claude-3.5-Sonnet | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| Claude-3.7-Sonnet | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| Gemini-2.0-Flash | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
| Gemini-2.5-Pro | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 | 0/30 |
A.12 Ablation on paraphrased prompt templates
In the natural suite, we use 5 paraphrased templates for problem statements and 30 for operations (e.g., “A space tunnel connects planet1 and planet2”) to smooth out ambiguity in natural language descriptions. However, in the main and challenge suites, we use formalized prompt templates to enable large-scale evaluation. We conducted an ablation to test how prompt template variations affect performance. Prompting GPT-4o with “Paraphrase the following description,” we generated five paraphrases and found that overall performance trends remained consistent.
| o4-mini | GPT-4.1 | |||
| Task | Default | Paraphrased | Default | Paraphrased |
| Priority Queue | 0.99 | 0.84 | 0.39 | 0.36 |
| Trie Tree | 0.89 | 0.90 | 0.63 | 0.61 |
| Geom Graph | 0.37 | 0.37 | 0.20 | 0.14 |
A.13 Preliminary study on strategy adaptation with DSR-Bench
In Section˜5, we note a future direction is to study whether models can select appropriate data structures given task requirements and switch dynamically. DSR-Bench already implicitly evaluates this adaptivity by not predefining strategies in prompts, requiring models to adjust their reasoning to the task. This emerges in (i) contrasting but related structure pairs (e.g., BFS/DFS, queue/priority queue), (ii) length generalization that demands adaptive complexity management, (iii) distribution shifts (Section˜4.2) requiring robustness across input patterns, and (iv) natural language scenarios that demand transferring strategies across contexts (Section˜4.3).
Beyond implicit evaluation, DSR-Bench’s modular design—data generation, prompt templates, automated evaluation, and schema-based verification—readily supports dynamic strategy switching. A preliminary study is shown in Table˜36 and Table˜35. We mask structure names or add constraints to DSR-Bench’s prompts, and test whether models can, (i) select BFS vs DFS for graph tasks, or (ii) choose between K-D tree and array for nearest-neighbor queries under varying complexities.
This extension underscores DSR-Bench’s broader value: it serves as a fundamental, extensible framework for probing not only structural reasoning but also more advanced reasoning abilities.
| Task | GPT-4.1 | o4-mini |
| Shortest Path | 9.5/10 (BFS) | 10/10 (BFS) |
| Cycle Detection | 10/10 (DFS) | 10/10 (DFS) |
| Connectivity Detection | 10/10 (Both equal; 8 BFS, 2 DFS) | 10/10 (Both equal; 6 BFS, 4 DFS) |
| Task | GPT-4.1 | o4-mini |
| 1 query, n=20–30 | 9/10 (array) | 10/10 (array) |
| 1 query, n=40–60 | 6/10 (array) | 10/10 (array) |
| n queries, n=5–10 | 10/10 (array) | 10/10 (array) |
| n queries, n=20–30 | 10/10 (kd-tree) | 10/10 (kd-tree) |